A sample of 4 different calculators is randomly selected from a group containing 16 that are defective and 31 that have no defects. What is the probability that at least one of the calculators is defective?
Answer by stanbon(75887)

A sample of 4 different calculators is randomly selected from a group containing 16 that are defective and 31 that have no defects. What is the probability that at least one of the calculators is defective?
-------
Binomial Problem with n = 4 ; p(defect) = 16/47 ; p(not) = 31/47
------
P(at least 1 defect) = 1 - P(none of 4 is defect)
----
= 1 - (31/47)^4
----
= 0.8107
====================
Chyeers,
Stan H.
====================